OpenStudy (anonymous):

Find the inverse if the following equation. Show work please. f(x)=2/3x+1

OpenStudy (anonymous):

to find the inverse of an equation you swap all your y's with x's, and all your x's with y's. You'll get a new equation and you need to solve for y/

OpenStudy (anonymous):

\[f(x) = 2/(3x+1)\] \[f^-1(x) = 2/(3y+1)\] \[3y +1 = 2/x\] \[3y = -1 + 2/x\] \[y = -1/3 + 2/3x\] \[f^-1(x) = -1/3 + 2/3x\]

OpenStudy (anonymous):

Thank you.